# Data curation, Methodology, Writing - original draft

### v3

@@ 3.4 Loss Function
Following previous methods [45][44][46], our model is trained end-to-end using the deep supervision strategy [50]. As illustrated in Fig. 2, final segmentation results are generated by the segmentation head ($1\times 1$ convolutional layer). Additionally, two smaller resolutions of decoder outputs are selected as auxiliary supervision signals. The deep supervision path in Fig. 2 consists of an upsample layer and a $1\times 1$ convolutional layer. Therefore, the loss function can be formulated as follows:
@@ 4.1 Datasets
To thoroughly compare MOSformer to previous methods, we conduct experiments on three challenging benchmarks: the Synapse multi-organ segmentation dataset [51], the automated cardiac diagnosis challenge (ACDC) dataset [52], and the abdominal organ segmentation (AMOS) dataset [53].
Synapse for Multi-organ Segmentation. This dataset consists of 30 abdominal CT scans with 8 organs (aorta, gallbladder, left kidney, right kidney, liver, pancreas, spleen, and stomach). Each volume has $85\sim 198$ slices of $512\times 512$ pixels. Following the splits adopted in TransUNet [23], the dataset is divided into 18 training cases and 12 testing cases.
ACDC for Automated Cardiac Diagnosis Challenge. The ACDC dataset includes cardiac MRI images of 100 patients from real clinical exams with manual annotations of left ventricle (LV), right ventricle (RV), and myocardium (Myo). Consistent with TransUNet [23], the dataset is split into 70 training cases, 10 validation cases, and 20 testing cases.
@@ 5.1 Comparisons with SOTAs
Table 1: Comparison with state-of-the-art models on the multi-organ segmentation (Synapse) dataset. The best results are highlighted in bold and the second-best results are underlined. The evaluation metrics are DSC and HD95, consistent with TransUNet [23]. Moreover, DSC of each organ is reported in this table. and indicate the results are borrowed from [46] and [22], respectively. means the baselines are implemented by ourselves. Baselines without any symbol represent the results are from the original papers.
@@
2D UNet [12] [MICCAI’15] 76.85 39.70 89.07 69.72 77.77 68.60 93.43 53.98 86.67 75.58
AttnUNet [54] [MedIA’19] 77.77 36.02 89.55 68.88 77.98 71.11 93.57 58.04 87.30 75.75
TransUNet [23] [MedIA’24] 77.48 31.69 87.23 63.13 81.87 77.02 94.08 55.86 85.08 75.62
MISSFormer [43] [TMI’23] 81.96 18.20 86.99 68.65 85.21 82.00 94.41 65.67 91.92 80.81
SwinUNet [22] [ECCVW’22] 79.12 21.55 85.47 66.53 83.28 79.61 94.29 56.58 90.66 76.60
MT-UNet [55] [ICASSP’22] 78.59 26.59 87.92 64.99 81.47 77.29 93.06 59.46 87.75 76.81
UCTransNet [56] [AAAI’22] 78.23 26.75 88.86 66.97 80.19 73.18 93.17 56.22 87.84 79.43
CASTformer [42] [NeurIPS’22] 82.55 22.73 89.05 67.48 86.05 82.17 95.61 67.49 91.00 81.55
HiFormer [57] [WACV’23] 80.39 14.70 86.21 65.69 85.23 79.77 94.61 59.52 90.99 81.08
SAMed [58] [arXiv’23] 81.88 20.64 87.77 69.11 80.45 79.95 94.80 72.17 88.72 82.06
3D V-Net [59] [3DV’16] 68.81 - 75.34 51.87 77.10 80.75 87.84 40.05 80.56 56.98
CoTr [45] [MICCAI’21] 80.78 19.15 85.42 68.93 85.45 83.62 93.89 63.77 88.58 76.23
UNETR [44] [WACV’22] 79.56 22.97 89.99 60.56 85.66 84.80 94.46 59.25 87.81 73.99
SwinUNETR [60] [MICCAIW’22] 83.51 14.78 90.75 66.72 86.51 85.88 95.33 70.07 94.59 78.20
nnFormer [46] [TIP’23] 86.57 10.63 92.04 70.17 86.57 86.25 96.84 83.35 90.51 86.83
SAM3D [61] [ISBI’24] 79.56 17.87 89.57 49.81 86.31 85.64 95.42 69.32 84.29 76.11
2.5D AFTer-UNet [27] [WACV’22] 81.02 - 90.91 64.81 87.90 85.30 92.20 63.54 90.99 72.48
TransUNet-2.5D [62] [TIM’23] 84.24 19.24 89.97 73.28 83.99 81.33 95.61 70.39 94.59 84.78
CSA-Net [38] [CIBM’24] 79.96 32.11 83.91 64.99 83.56 79.93 94.43 62.65 91.12 79.12
@@
We select several state-of-the-art 2D, 3D, and 2.5D medical image segmentation models as our baselines. To ensure a fair comparison, all models are trained and evaluated using identical preprocessing pipelines. Specifically, we apply the TransUNet preprocessing protocol [23] to both the multi-organ segmentation (Synapse) and the automated cardiac diagnosis challenge (ACDC) datasets, and follow the preprocessing procedures of [53] for the abdominal organ segmentation (AMOS) dataset. It should be noted that we only visualize selected qualitative results from some representative models for clarity and visual impact.
Multi-organ Segmentation (Synapse). Quantitative results of state-of-the-art models and our MOSformer are presented in Table 1. MOSformer achieves $85.63\%$ DSC and $13.40$ $\rm mm$} HD95 on this dataset. Compared with the best 2D-based method, i.e., CASTformer [42], MOSformer is able to surpass it by a large margin ($3.08\%$ DSC and 9.33-9.33 $\rm mm$} HD95). For 2.5D-based baselines, MOSformer demonstrates notable performance enhancements, offering at least $4.61\%$, $1.39\%$, and $5.67\%$ DSC gains over AFTer-UNet [27], TransUNet-2.5D [62], and CSA-Net [38], respectively. These results indicate i) the necessity of inter-slice information in 3D medical image segmentation; and ii) the effectiveness of distinguishable and consistent slice features produced by dual encoders with a momentum update.
We also compare our MOSformer with 3D-based segmentation methods. It still has competitive performance, surpassing five of the most widely recognized models and achieving comparable performance to nnFormer [46]. It should be noted that MOSformer obtains better DSC than nnFormer in four organs (half of the categories), including gallbladder ($1.73\%$), left kidney ($3.75\%$), spleen ($1.78\%$), and stomach ($1.04\%$). Among these organs, gallbladder and stomach are two of the most difficult organs to segment since the gallbladder is very small and the boundaries between the gallbladder and the liver are blurred while the stomach has a significant intra-class variance. This reveals that our MOSformer can learn more discriminative features and has a comprehensive understanding of organ structures.
@@
Table 2: Comparison with the state-of-the-art models on the automated cardiac diagnosis challenge (ACDC) dataset. The best results are highlighted in bold and the second-best results are underlined. We only report DSC in this table, following the evaluation setting of TransUNet [23]. Moreover, DSC of each anatomical structure is reported in this table. and indicate the results are borrowed from [46] and [55], respectively. means the baselines are implemented by ourselves. Baselines without any symbol represent the results are from the original papers.
@@
2D UNet [12] [MICCAI’15] 87.60 84.62 84.52 93.68
AttnUNet [54] [MedIA’19] 86.90 83.27 84.33 93.53
TransUNet [23] [MedIA’24] 89.71 86.67 87.27 95.18
MISSFormer [43] [TMI’23] 91.19 89.85 88.38 95.34
SwinUNet [22] [ECCVW’22] 88.07 85.77 84.42 94.03
MT-UNet [55] [ICASSP’22] 90.43 86.64 89.04 95.62
UCTransNet [56] [AAAI’22] 91.98 90.06 89.87 96.02
HiFormer [57] [WACV’23] 90.40 88.24 87.63 95.30
3D UNETR [44] [WACV’22] 88.61 85.29 86.52 94.02
nnFormer [46] [TIP’23] 92.06 90.94 89.58 95.65
SAM3D [61] [ISBI’24] 90.41 89.44 87.12 94.67
2.5D CAT-Net [28] [TMI’22] 90.02 86.05 88.75 95.27
CSA-Net [38] [CIBM’24] 89.58 86.56 86.91 95.26
@@
Automated Cardiac Diagnosis Challenge (ACDC). To further prove the model’s generalization performance, MOSformer is evaluated on the automated cardiac diagnosis challenge (ACDC) dataset. It should be noted that MRI images in this dataset can be considered anisotropic since they have high in-plane image resolution (e.g., $1.37\sim 1.68$ $\rm mm$}) and low through-plane resolution (e.g., 55 $\rm mm$}[52]. Quantitative results are summarized in Table 2. Compared with state-of-the-art methods (2D, 2.5D, and 3D-based), MOSformer achieves the best performance with $92.19\%$ DSC. Thus, the above results indicate that our 2.5D-based MOSformer is more effective at processing anisotropic data compared with 3D-based models. Fig. 5 presents qualitative comparisons for different methods on this dataset. As seen, MOSformer can locate anatomical structures more accurately. Specifically, in case 3, many models mistakenly classify regions outside the myocardium into the right ventricle while MOSformer does not produce any false positive predictions.
@@
2D UNet [12] [MICCAI’15] 82.53 92.25 92.45 92.50 81.85 79.98 94.73 84.80 92.20 82.94 77.35 67.13 69.34 72.77 82.40 75.31
TransUNet [23] [MedIA’24] 80.10 91.26 92.47 91.90 78.01 77.00 94.93 80.04 91.98 82.99 74.30 63.66 53.84 71.65 81.37 76.03
MISSFormer [43] [TMI’23] 78.16 93.13 91.98 91.88 75.89 71.87 94.27 80.14 88.74 77.53 71.39 60.65 59.32 64.43 77.97 73.16
UCTransNet [56] [AAAI’22] 82.34 93.37 92.32 91.90 77.09 79.77 94.78 85.95 91.77 82.84 77.44 65.88 68.98 71.36 83.93 77.71
HiFormer [57] [WACV’23] 80.03 92.73 92.79 92.01 79.44 76.42 94.55 82.65 90.56 80.16 73.59 61.14 58.73 68.12 82.01 75.64
3D UNETR [44] [WACV’22] 78.07 93.38 93.00 92.28 73.17 69.72 94.86 73.25 90.82 80.20 73.44 65.19 60.69 65.46 74.10 71.49
nnFormer [46] [TIP’23] 78.66 91.43 92.39 92.08 76.74 69.16 94.95 84.84 89.53 82.06 75.91 62.56 60.36 68.50 74.74 64.61
2.5D CSA-Net [38] [CIBM’24] 82.12 91.25 93.51 93.68 79.01 78.80 95.32 82.14 91.64 83.94 75.18 68.27 69.37 71.36 83.00 75.33
@@
MISSFormer [43] [TMI’23] 81.96 < 0.001 UCTransNet [56] [AAAI’22] 91.89 0.745 UNet [12] [MICCAI’15] 82.53 < 0.001
nnFormer [46] [TIP’23] 86.57 0.339 nnFormer [46] [TIP’23] 92.06 0.826 nnFormer [46] [TIP’23] 78.66 < 0.001
CSA-Net [38] [CIBM’24] 79.96 < 0.001 CAT-Net [28] [TMI’22] 90.02 < 0.001 CSA-Net [38] [CIBM’24] 82.12 < 0.001
@@
Figure 7: Visualization of embedding space learned under three encoder settings on the multi-organ segmentation (Synapse) test set (1,568 slices). Each point represents the feature of a slice. Distinct colors are used to differentiate embeddings from different encoders. Because Model-2 uses a single encoder for both target and neighborhood slices, their embeddings are identical. Therefore, only the blue points are shown. Dimensions are reduced by t-SNE [63]. (a) Model-2 (Single encoder); (b) Model-3 (Dual encoders updated independently); (c) MOSformer (Dual encoders with a momentum update).
@@ 5.2 Ablation Study
Furthermore, we also adopt t-SNE [63] to visualize the encoded embedding space learned from three encoder settings on the multi-organ segmentation (Synapse) test set. Model-2 employs a single encoder to process both target and neighborhood slices. Consequently, target and neighborhood slice features originate from the same feature space, as depicted in Fig. 7 (a). This setup poses challenges for the model in distinguishing individual slices and acquiring slice-specific information during inter-slice fusion. In contrast, the embedding space learned by dual encoders is distinguishable, as illustrated in Fig. 7 (b) and (c). It can also be observed that incorporating the momentum update in dual encoders facilitates consistency among slice features, as shown in Fig. 7 (c), thereby further boosting segmentation performance.
Figure 8: Class activation maps of the gallbladder and the stomach categories (from top to bottom) produced by Grad-CAM [64]. The class activation maps are generated from the last decoder layer.
@@
Efficacy of The Inter-slice Fusion Transformer. We remove IF-Trans modules in MOSformer and obtain a baseline model (Model-1). Compared with the baseline model, models incorporating IF-Trans (Model-2 and MOSformer) offer substantial improvements on the Synapse ($1.81\%$ and $3.21\%$ DSC gains), the ACDC dataset ($0.43\%$ and $0.58\%$ DSC gains), and the AMOS dataset ($1.35\%$ and $4.15\%$ DSC gains). These results demonstrate that inter-slice information is crucial in 3D medical image segmentation and the proposed IF-Trans can effectively fuse such information.
Additionally, we employ Grad-CAM [64] to visualize discriminative regions of the models, as depicted in Fig. 8. Compared with baseline Model-1, we can see that inter-slice information is beneficial, but Model-2 and Model-3 still tend to assign weights to irrelevant regions. Distinguishable and consistent inter-slice features within MOSformer can address the above issue, demonstrating enhanced precision in localizing organs of interest.

### v4

@@ 3.4 Loss Function
Following previous methods [xie2021cotr, hatamizadeh2022unetr, zhou2023nnformer], our model is trained end-to-end using the deep supervision strategy [lee2015deeply]. As illustrated in Fig. 2, final segmentation results are generated by the segmentation head ($1\times 1$ convolutional layer). Additionally, two smaller resolutions of decoder outputs are selected as auxiliary supervision signals. The deep supervision path in Fig. 2 consists of an upsample layer and a $1\times 1$ convolutional layer. Therefore, the loss function can be formulated as follows:
@@ 4.1 Datasets
To thoroughly compare MOSformer to previous methods, we conduct experiments on three challenging benchmarks: the Synapse multi-organ segmentation dataset [landman2015miccai], the automated cardiac diagnosis challenge (ACDC) dataset [bernard2018deep], and the abdominal organ segmentation (AMOS) dataset [ji2022amos].
Synapse for Multi-organ Segmentation. This dataset consists of 30 abdominal CT scans with 8 organs (aorta, gallbladder, left kidney, right kidney, liver, pancreas, spleen, and stomach). Each volume has $85\sim 198$ slices of $512\times 512$ pixels. Following the splits adopted in TransUNet [chen2024transunet], the dataset is divided into 18 training cases and 12 testing cases.
ACDC for Automated Cardiac Diagnosis Challenge. The ACDC dataset includes cardiac MRI images of 100 patients from real clinical exams with manual annotations of left ventricle (LV), right ventricle (RV), and myocardium (Myo). Consistent with TransUNet [chen2024transunet], the dataset is split into 70 training cases, 10 validation cases, and 20 testing cases.
@@ 5.1 Comparisons with SOTAs
Table 1: Comparison with state-of-the-art models on the multi-organ segmentation (Synapse) dataset. The best results are highlighted in bold and the second-best results are underlined. The evaluation metrics are DSC and HD95, consistent with TransUNet [chen2024transunet]. Moreover, DSC of each organ is reported in this table. and indicate the results are borrowed from [zhou2023nnformer] and [cao2022swin], respectively. means the baselines are implemented by ourselves. Baselines without any symbol represent the results are from the original papers.
@@
2D UNet [ronneberger2015u] [MICCAI’15] 76.85 39.70 89.07 69.72 77.77 68.60 93.43 53.98 86.67 75.58
AttnUNet [schlemper2019attention] [MedIA’19] 77.77 36.02 89.55 68.88 77.98 71.11 93.57 58.04 87.30 75.75
TransUNet [chen2024transunet] [MedIA’24] 77.48 31.69 87.23 63.13 81.87 77.02 94.08 55.86 85.08 75.62
MISSFormer [huang2023missformer] [TMI’23] 81.96 18.20 86.99 68.65 85.21 82.00 94.41 65.67 91.92 80.81
SwinUNet [cao2022swin] [ECCVW’22] 79.12 21.55 85.47 66.53 83.28 79.61 94.29 56.58 90.66 76.60
MT-UNet [wang2022mixed] [ICASSP’22] 78.59 26.59 87.92 64.99 81.47 77.29 93.06 59.46 87.75 76.81
UCTransNet [wang2022uctransnet] [AAAI’22] 78.23 26.75 88.86 66.97 80.19 73.18 93.17 56.22 87.84 79.43
CASTformer [you2022class] [NeurIPS’22] 82.55 22.73 89.05 67.48 86.05 82.17 95.61 67.49 91.00 81.55
HiFormer [heidari2023hiformer] [WACV’23] 80.39 14.70 86.21 65.69 85.23 79.77 94.61 59.52 90.99 81.08
SAMed [zhang2023customized] [arXiv’23] 81.88 20.64 87.77 69.11 80.45 79.95 94.80 72.17 88.72 82.06
3D V-Net [milletari2016v] [3DV’16] 68.81 - 75.34 51.87 77.10 80.75 87.84 40.05 80.56 56.98
CoTr [xie2021cotr] [MICCAI’21] 80.78 19.15 85.42 68.93 85.45 83.62 93.89 63.77 88.58 76.23
UNETR [hatamizadeh2022unetr] [WACV’22] 79.56 22.97 89.99 60.56 85.66 84.80 94.46 59.25 87.81 73.99
SwinUNETR [hatamizadeh2021swin] [MICCAIW’22] 83.51 14.78 90.75 66.72 86.51 85.88 95.33 70.07 94.59 78.20
nnFormer [zhou2023nnformer] [TIP’23] 86.57 10.63 92.04 70.17 86.57 86.25 96.84 83.35 90.51 86.83
SAM3D [bui2024sam3d] [ISBI’24] 79.56 17.87 89.57 49.81 86.31 85.64 95.42 69.32 84.29 76.11
2.5D AFTer-UNet [yan2022after] [WACV’22] 81.02 - 90.91 64.81 87.90 85.30 92.20 63.54 90.99 72.48
TransUNet-2.5D [zhang2023multi] [TIM’23] 84.24 19.24 89.97 73.28 83.99 81.33 95.61 70.39 94.59 84.78
CSA-Net [kumar2024flexible] [CIBM’24] 79.96 32.11 83.91 64.99 83.56 79.93 94.43 62.65 91.12 79.12
@@
We select several state-of-the-art 2D, 3D, and 2.5D medical image segmentation models as our baselines. To ensure a fair comparison, all models are trained and evaluated using identical preprocessing pipelines. Specifically, we apply the TransUNet preprocessing protocol [chen2024transunet] to both the multi-organ segmentation (Synapse) and the automated cardiac diagnosis challenge (ACDC) datasets, and follow the preprocessing procedures of [ji2022amos] for the abdominal organ segmentation (AMOS) dataset. It should be noted that we only visualize selected qualitative results from some representative models for clarity and visual impact.
Multi-organ Segmentation (Synapse). Quantitative results of state-of-the-art models and our MOSformer are presented in Table 1. MOSformer achieves $85.63\%$ DSC and $13.40$ $\rm mm$} HD95 on this dataset. Compared with the best 2D-based method, i.e., CASTformer [you2022class], MOSformer is able to surpass it by a large margin ($3.08\%$ DSC and 9.33-9.33 $\rm mm$} HD95). For 2.5D-based baselines, MOSformer demonstrates notable performance enhancements, offering at least $4.61\%$, $1.39\%$, and $5.67\%$ DSC gains over AFTer-UNet [yan2022after], TransUNet-2.5D [zhang2023multi], and CSA-Net [kumar2024flexible], respectively. These results indicate i) the necessity of inter-slice information in 3D medical image segmentation; and ii) the effectiveness of distinguishable and consistent slice features produced by dual encoders with a momentum update.
@@
Table 2: Comparison with the state-of-the-art models on the automated cardiac diagnosis challenge (ACDC) dataset. The best results are highlighted in bold and the second-best results are underlined. We only report DSC in this table, following the evaluation setting of TransUNet [chen2024transunet]. Moreover, DSC of each anatomical structure is reported in this table. and indicate the results are borrowed from [zhou2023nnformer] and [wang2022mixed], respectively. means the baselines are implemented by ourselves. Baselines without any symbol represent the results are from the original papers.
@@
2D UNet [ronneberger2015u] [MICCAI’15] 87.60 84.62 84.52 93.68
AttnUNet [schlemper2019attention] [MedIA’19] 86.90 83.27 84.33 93.53
TransUNet [chen2024transunet] [MedIA’24] 89.71 86.67 87.27 95.18
MISSFormer [huang2023missformer] [TMI’23] 91.19 89.85 88.38 95.34
SwinUNet [cao2022swin] [ECCVW’22] 88.07 85.77 84.42 94.03
MT-UNet [wang2022mixed] [ICASSP’22] 90.43 86.64 89.04 95.62
UCTransNet [wang2022uctransnet] [AAAI’22] 91.98 90.06 89.87 96.02
HiFormer [heidari2023hiformer] [WACV’23] 90.40 88.24 87.63 95.30
3D UNETR [hatamizadeh2022unetr] [WACV’22] 88.61 85.29 86.52 94.02
nnFormer [zhou2023nnformer] [TIP’23] 92.06 90.94 89.58 95.65
SAM3D [bui2024sam3d] [ISBI’24] 90.41 89.44 87.12 94.67
2.5D CAT-Net [hung2022cat] [TMI’22] 90.02 86.05 88.75 95.27
CSA-Net [kumar2024flexible] [CIBM’24] 89.58 86.56 86.91 95.26
@@
We also compare our MOSformer with 3D-based segmentation methods. It still has competitive performance, surpassing five of the most widely recognized models and achieving comparable performance to nnFormer [zhou2023nnformer]. It should be noted that MOSformer obtains better DSC than nnFormer in four organs (half of the categories), including gallbladder ($1.73\%$), left kidney ($3.75\%$), spleen ($1.78\%$), and stomach ($1.04\%$). Among these organs, gallbladder and stomach are two of the most difficult organs to segment since the gallbladder is very small and the boundaries between the gallbladder and the liver are blurred while the stomach has a significant intra-class variance. This reveals that our MOSformer can learn more discriminative features and has a comprehensive understanding of organ structures.
Automated Cardiac Diagnosis Challenge (ACDC). To further prove the model’s generalization performance, MOSformer is evaluated on the automated cardiac diagnosis challenge (ACDC) dataset. It should be noted that MRI images in this dataset can be considered anisotropic since they have high in-plane image resolution (e.g., $1.37\sim 1.68$ $\rm mm$}) and low through-plane resolution (e.g., 55 $\rm mm$}[bernard2018deep]. Quantitative results are summarized in Table 2. Compared with state-of-the-art methods (2D, 2.5D, and 3D-based), MOSformer achieves the best performance with $92.19\%$ DSC. Thus, the above results indicate that our 2.5D-based MOSformer is more effective at processing anisotropic data compared with 3D-based models. Fig. 5 presents qualitative comparisons for different methods on this dataset. As seen, MOSformer can locate anatomical structures more accurately. Specifically, in case 3, many models mistakenly classify regions outside the myocardium into the right ventricle while MOSformer does not produce any false positive predictions.
@@
2D UNet [ronneberger2015u] [MICCAI’15] 82.53 92.25 92.45 92.50 81.85 79.98 94.73 84.80 92.20 82.94 77.35 67.13 69.34 72.77 82.40 75.31
TransUNet [chen2024transunet] [MedIA’24] 80.10 91.26 92.47 91.90 78.01 77.00 94.93 80.04 91.98 82.99 74.30 63.66 53.84 71.65 81.37 76.03
MISSFormer [huang2023missformer] [TMI’23] 78.16 93.13 91.98 91.88 75.89 71.87 94.27 80.14 88.74 77.53 71.39 60.65 59.32 64.43 77.97 73.16
UCTransNet [wang2022uctransnet] [AAAI’22] 82.34 93.37 92.32 91.90 77.09 79.77 94.78 85.95 91.77 82.84 77.44 65.88 68.98 71.36 83.93 77.71
HiFormer [heidari2023hiformer] [WACV’23] 80.03 92.73 92.79 92.01 79.44 76.42 94.55 82.65 90.56 80.16 73.59 61.14 58.73 68.12 82.01 75.64
3D UNETR [hatamizadeh2022unetr] [WACV’22] 78.07 93.38 93.00 92.28 73.17 69.72 94.86 73.25 90.82 80.20 73.44 65.19 60.69 65.46 74.10 71.49
nnFormer [zhou2023nnformer] [TIP’23] 78.66 91.43 92.39 92.08 76.74 69.16 94.95 84.84 89.53 82.06 75.91 62.56 60.36 68.50 74.74 64.61
2.5D CSA-Net [kumar2024flexible] [CIBM’24] 82.12 91.25 93.51 93.68 79.01 78.80 95.32 82.14 91.64 83.94 75.18 68.27 69.37 71.36 83.00 75.33
@@
MISSFormer [huang2023missformer] [TMI’23] 81.96 < 0.001 UCTransNet [wang2022uctransnet] [AAAI’22] 91.89 0.745 UNet [ronneberger2015u] [MICCAI’15] 82.53 < 0.001
nnFormer [zhou2023nnformer] [TIP’23] 86.57 0.339 nnFormer [zhou2023nnformer] [TIP’23] 92.06 0.826 nnFormer [zhou2023nnformer] [TIP’23] 78.66 < 0.001
CSA-Net [kumar2024flexible] [CIBM’24] 79.96 < 0.001 CAT-Net [hung2022cat] [TMI’22] 90.02 < 0.001 CSA-Net [kumar2024flexible] [CIBM’24] 82.12 < 0.001
@@
Figure 7: Visualization of embedding space learned under three encoder settings on the multi-organ segmentation (Synapse) test set (1,568 slices). Each point represents the feature of a slice. Distinct colors are used to differentiate embeddings from different encoders. Because Model-2 uses a single encoder for both target and neighborhood slices, their embeddings are identical. Therefore, only the blue points are shown. Dimensions are reduced by t-SNE [van2008visualizing]. (a) Model-2 (Single encoder); (b) Model-3 (Dual encoders updated independently); (c) MOSformer (Dual encoders with a momentum update).
@@ 5.2 Ablation Study
Furthermore, we also adopt t-SNE [van2008visualizing] to visualize the encoded embedding space learned from three encoder settings on the multi-organ segmentation (Synapse) test set. Model-2 employs a single encoder to process both target and neighborhood slices. Consequently, target and neighborhood slice features originate from the same feature space, as depicted in Fig. 7 (a). This setup poses challenges for the model in distinguishing individual slices and acquiring slice-specific information during inter-slice fusion. In contrast, the embedding space learned by dual encoders is distinguishable, as illustrated in Fig. 7 (b) and (c). It can also be observed that incorporating the momentum update in dual encoders facilitates consistency among slice features, as shown in Fig. 7 (c), thereby further boosting segmentation performance.
Figure 8: Class activation maps of the gallbladder and the stomach categories (from top to bottom) produced by Grad-CAM [selvaraju2017grad]. The class activation maps are generated from the last decoder layer.
@@
Efficacy of The Inter-slice Fusion Transformer. Compared to the baseline Model-1, Model-2 with the IF-Trans module offers substantial improvements, increasing DSC by $1.81\%$, $0.43\%$, and $1.35\%$ on the Synapse, ACDC, and AMOS datasets, respectively. Furthermore, enhancing feature discriminability with dual encoders yields even greater performance gains. Specifically, Model-3 and our MOSformer further boost DSC by $2.51\%$ and $3.21\%$ on Synapse, $0.49\%$ and $0.58\%$ on ACDC, and $2.60\%$ and $4.15\%$ on AMOS, respectively.
Additionally, we employ Grad-CAM [selvaraju2017grad] to visualize discriminative regions of the models, as depicted in Fig. 8. Compared with baseline Model-1, we can see that inter-slice information is beneficial, but Model-2 and Model-3 still tend to assign weights to irrelevant regions. Distinguishable and consistent inter-slice features within MOSformer can address the above issue, demonstrating enhanced precision in localizing organs of interest.